Four times the measure of the angle is 25 more than one third the measure of its supplement
Answer by checkley77(12569)

4x=25+1/3(180-x)
4x=25+60-x/3
4x+x/3=85
(4x*3+x)/3=85
(12x+x)/3=85
13x=3*85
13x=255
x=255/13
x=19.6154 ans.
4*19.6154=25+1/3(180-19.6154)
78.4616=25+1/3*160.3846
78.4616=25+52.4615
78.4616~78.4615